OpenStudy (anonymous):

Solve P = 2w + 2h for w. w = P – 2h w = P – 2h – 2 w = P-h/2 w = P-2h/2

OpenStudy (diyadiya):

First subtract 2h from both sides can you do that?

OpenStudy (anonymous):

Yes

OpenStudy (diyadiya):

What do you get?

OpenStudy (diyadiya):

What ? subtract 2h from both sides P-2h = 2w +2h - 2h

OpenStudy (diyadiya):

Now?

OpenStudy (diyadiya):

P - 2h = 2w Now divide both sides by 2 what do yu get on both sides?

OpenStudy (anonymous):

P = 2w + 2h P = 2(w + h) P/2 = w + h P/2 - h = w w = P/2 - h :o)

OpenStudy (anonymous):

Is It Correct?

OpenStudy (diyadiya):

Yes that is correct But your choices doesn't have that lets try another easier way P = 2w + 2h P-2h =2w divide both sides by 2 (P-2h)/2 = w got it?

OpenStudy (anonymous):

Yes.

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

C is the answer?

OpenStudy (diyadiya):

how?

OpenStudy (diyadiya):

w = P-2h/2 this is what we got

OpenStudy (anonymous):

Or D?

OpenStudy (anonymous):

Correct

OpenStudy (diyadiya):

Yes!

OpenStudy (anonymous):

That Is What We Got.

OpenStudy (diyadiya):

Yes

OpenStudy (anonymous):

Yay! Thanks!

OpenStudy (diyadiya):

No problem :)